OpenStudy (haleyelizabeth2017):

Solve by elimination~please show all work! -2x+2y+3z=0 -2x-y+z=-3 2x+3y+3z=5 I think it is no solution because when I solved for it, x=1.75, y=0, z=.5 I plugged it back into the equations and it was good for the second two, but it was wrong for the first one. Can someone please show steps to see if I did it right?

OpenStudy (haleyelizabeth2017):

@SydtheKid913 @Blu-Girl @TheSmartOne please help! :) thanks!

myininaya (myininaya):

So I would try subtracting the first two equations together first to have an equation with just y and z. I could get another equation just involving y and z by adding the last two equations. So then you could solve that 2-variable system for y and z.

myininaya (myininaya):

And then go back to one of your initial equations to find x.

myininaya (myininaya):

but you should see something kinda neat when you are trying to solve for y and z that will kinda give you a little more confidence about the conclusion you have already reached.

OpenStudy (haleyelizabeth2017):

I'm still really confused :(

OpenStudy (haleyelizabeth2017):

Thats what I did. One second and I will show you the steps I did @myininaya :)

OpenStudy (haleyelizabeth2017):

This is what I did :/ sorry it's sloppy @myininaya :(

myininaya (myininaya):

I think I actually made a mistake Your first solution might be okay. I will check it.

OpenStudy (haleyelizabeth2017):

Thank you so much :)

OpenStudy (haleyelizabeth2017):

I think I made the mistake. I plugged it in the calculator thingy and it should be x=1 y=1 z=0

myininaya (myininaya):

\[3y+2z=3 \\ 2y+4z=2 \\ \text{ rewrite second equation as } y+2z=1 \\ \text{ so we have } \\ 3y+2z=3 \\ -(y+2z=1) \\2y=2 \\ y=1 \] Trying to look at your work and see how you got y=1/2 ...

myininaya (myininaya):

Oh I see what you did

myininaya (myininaya):

You just change that one number from 2 to 6

myininaya (myininaya):

you had 2y+4z=2 but you wrote 2y+4z=6

myininaya (myininaya):

when you were taking away from that other equation

myininaya (myininaya):

that is how you got y=0

myininaya (myininaya):

so that step there should have looked more like this: 6y+4y=6 -(2y+4z=2) ------- 4y=4 y=1

myininaya (myininaya):

Then you would have a difference z and therefore a different x

OpenStudy (haleyelizabeth2017):

Okay. thank you so much! I really appreciate it! I might need some more help, but sublimation instead for a different problem :)

myininaya (myininaya):

k

OpenStudy (haleyelizabeth2017):

it might be a sec because I have to rewrite all of my work :)

myininaya (myininaya):

Do you see what I was talking about though? Or understand where the error occurred?

OpenStudy (haleyelizabeth2017):

yes. thank you again!

OpenStudy (haleyelizabeth2017):

but then z=1....

OpenStudy (haleyelizabeth2017):

I found what I did....again

myininaya (myininaya):

z isn't 1

OpenStudy (haleyelizabeth2017):

its 0. I accidently put 2-2 as 4 :P haha I'm too tired to do math :/

myininaya (myininaya):

lol

myininaya (myininaya):

we all are sometimes

OpenStudy (haleyelizabeth2017):

haha but now I have to rewrite it, yet again :'(

OpenStudy (haleyelizabeth2017):

I'm switching to pencil now so I don't have to start over when I make a mistake :P

myininaya (myininaya):

I didn't correct your check section.

OpenStudy (haleyelizabeth2017):

It's okay. I got it now.....one second. I am starting the sublimation one :)

OpenStudy (haleyelizabeth2017):

Okay. I have the problem now :)